OpenStudy (anonymous):

Given the equation y − 3 = one half(x + 6) in point-slope form, identify the equation of the same line in standard form. x − 2y = −12 y = one halfx y = one halfx + 6 y = one halfx + 9 I think that the answer is option 1, but I want someone else to check please :)

OpenStudy (anonymous):

can you simplify the equation given to you? y − 3 = one half(x + 6) is really y − 3 = (1/2)(x + 6)

OpenStudy (anonymous):

Wouldn't it be y = 1/2x + 6 simplified?

OpenStudy (anonymous):

yep good job. do you see which solution it is now?

OpenStudy (anonymous):

Yeah I do now, but isn't standard form Ax + By = C

OpenStudy (anonymous):

i'd give you another medal if I could. you are right again. can you get the simplified eqatuion into stanard form?

OpenStudy (anonymous):

No, I don't think so

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

we have: y = 1/2x + 6 move the 6 to the other side, and move the y to the other side

OpenStudy (anonymous):

Would it be 6 = 1/2x + y?

OpenStudy (anonymous):

-6 = (1/2)x - y now times both sides by 2 to get rid of the fraction

OpenStudy (anonymous):

Is it -12 = x - 2y?

OpenStudy (anonymous):

you got it, do you see the answer up there to pick it?

OpenStudy (anonymous):

Answer 1?

OpenStudy (anonymous):

yep ^_^

OpenStudy (anonymous):

Thanks a lot!!!! :)

OpenStudy (anonymous):

Can you check this one for me too? I think the answer is option 4, I want to be positive. :) Given the equation y − 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in slope-intercept form. −three fourthsx + y = 10 3x − 4y = −40 y = three fourthsx + 12 y = three fourthsx + 10

OpenStudy (anonymous):

option 4 is what I come up with also ^_^